In a box,there are 3 red marbles,4 green marbles and 5 yellow marbles.What is the probability of drawing three marbles if draw simultaneously and
1.)they are of the same color?
2.) they are of 3 different colors?

Answer by Boreal(15235)

same color
if red, 3/12*2/11*1/10=6/1320=1/220
if green, 4*3*2 That =24/1320=4/220
if yellow, 5*4*3=60 That =60/1320=10/220
total is 15/220=3/44=0.068
==============
3 different colors
if red first, 3/12*4/9*5/8=60/864=5/72
if green first, 4/12*3/8*5/7=60/672=5/56
if yellow first, 5/12*3/7*4/6=60/504=5/42
=47040/169344=0.278